# Mojito: Motion Trajectory and Intensity Control for Video Generation

### v1

@@ Abstract
Recent advancements in diffusion models have shown great promise in producing high-quality video content. However, efficiently training diffusion models capable of integrating directional guidance and controllable motion intensity remains a challenging and under-explored area. This paper introduces Mojito, a diffusion model that incorporates both Motion trajectory and intensity control for text to video generation. Specifically, Mojito features a Directional Motion Control module that leverages cross-attention to efficiently direct the generated object’s motion without additional training, alongside a Motion Intensity Modulator that uses optical flow maps generated from videos to guide varying levels of motion intensity. Extensive experiments demonstrate Mojito’s effectiveness in achieving precise trajectory and intensity control with high computational efficiency, generating motion patterns that closely match specified directions and intensities, providing realistic dynamics that align well with natural motion in real-world scenarios.
@@ 1 Introduction
Video generation, and specifically text-to-video (T2V) generation [1, 14, 52, 23, 45, 47, 40], aims to produce diverse, high-quality videos that match given text prompts. Unlike image generation [34, 37, 28, 31, 49, 3], which focuses on generating a single static frame, video generation must ensure consistent and fluid motion across a sequence of frames. However, existing models struggle to provide flexible, controllable motion effects in response to user prompts, particularly when handling specific motion directions or varying levels of motion intensity. Moreover, these models often fall short in generating videos with large motion amplitudes or adaptable motion strengths that align with user intent. Figure LABEL:Fig:teaser illustrates these challenges, which Mojito effectively addresses through both an efficient, training-free method for direction control and a novel training-based approach for motion intensity modulation.
@@
Achieving precise control over both motion direction and intensity in video generation poses great challenges. First, capturing relative motion data is inherently complex in real-world videos due to the simultaneous movements of both cameras and objects; Second, there is a lack of large-scale, annotated datasets specifically for motion direction and intensity. Existing video datasets rarely include detailed labels for motion dynamics, and obtaining human annotations for these subtle aspects is both costly and labor-intensive; Third, training models with such detailed annotations would demand substantial computational resources.
@@
To address these challenges, we present Mojito, the first diffusion model for text-to-video generation that simultaneously integrates text-based prompts with flexible motion controls, allowing for the precise modulation of both motion direction and intensity. To achieve this, Mojito introduces two novel modules on top of the trained backbone: (1) the Directional Motion Control (DMC) module, which adjusts motion direction through cross-attention, aligning object trajectories with specified paths; and (2) the Motion Intensity Modulator (MIM) module, which encodes arbitrary motion intensities as features and integrates them seamlessly into the diffusion framework. Additionally, we explore an alternative by training a global motion intensity embedding layer as a conditional input to further enhance control over motion intensity.
@@
  • We introduce Mojito, a diffusion-based transformer model for text-to-video generation with integrated trajectory/direction guidance and motion intensity control.
  • We introduce a training-free, test-time directional guidance method that enables text-to-video generation with specified trajectories or directions.
  • We develop a training-based method for motion-guided generation that allows precise control over motion intensity to match user-defined motion levels.
  • We demonstrate that Mojito not only generates high-quality video content comparable to state-of-the-art models but also provides effective and efficient motion control. Through ablations, we provide insights into designing efficient, motion-controllable video generation techniques, emphasizing essential factors for advancing future motion-augmented video generative models.
@@ 2.1 Diffusion-based Text to Video Generation
With the rapid development of generative models, particularly diffusion models, numerous breakthroughs have been achieved in fields such as image generation [34] and video generation [14, 45, 52, 23, 5, 47, 40]. Regarding text-to-video (T2V) models, Make-A-Video [38], Imagen Video [36] are cascaded models, while most other works, such as LVDM [14], ModelScope [41], and Align your Latents [2], are Stable Diffusion-based models. They extend the SD framework to videos by incorporating temporal layers to ensure temporal consistency among frames and the spatial parameters are inherited from the pretrained SD UNet. Our model follows a similar architectural approach, using spatial transformers for spatial information processing and accepting directional control, and incorporating temporal transformers to manage temporal coherence.
@@ 2.2 Controllable Text-to-Visual Generation
Recent advancements in text-to-visual generation have enabled more precise and stable control over generated outputs. In image generation, several models [27, 17, 51, 29] have been developed to enhance control by leveraging semantic guidance from text prompts alongside various structural conditions. Some approaches utilize attention maps and bounding boxes to manage image layouts, enabling generation based on regional specifications [4, 46, 24, 6, 11, 13].  Mo et al. [26] presents a training-free method for controllable image generation; In video generation, which involves additional complexity due to motion, recent works have explored methods for motion control. Training-based methods, such as those in [50, 43], allow for trajectory-based motion control, while [19] achieves video control through designed masking in the diffusion process. Unlike these approaches, our method leverages attention maps in a training-free manner for directional motion control. Our model can control the movement of specific objects within the generated video, and uniquely, it enables motion intensity control—providing a new level of flexibility in video generation not addressed in previous works.
Figure 1: Overview of the Mojito framework. In the training pipeline (top),  Mojito uses a VAE Encoder to transform input frames into latent features, processed by Spatial and Temporal Transformers. Motion control is introduced through the Motion Intensity Encoder. During inference (bottom), Mojito generates videos following user-defined motion intensity and directional guidance. The Directional Motion Control Module interprets object phrases within the prompt to align attention with specified trajectories, enabling controlled motion dynamics in the generated video.
@@ 3 Preliminaries
Latent Diffusion Models (LDMs), such as Stable Diffusion [34], consist of two main components: an autoencoder and a denoising network. The autoencoder encodes images into a latent space, while the denoising network operates within this latent space, progressively transforming random noise into a coherent image representation. Building on this, Latent Video Diffusion Models (LVDMs) [14] extend LDMs to video generation by incorporating temporal dynamics in the latent space. They are trained using a noise-prediction objective, minimizing the difference between predicted and actual noise added to the data:
@@
4 Motion-Controlled Video Generation
Figure 1 illustrates the architecture of Mojito, which comprises two core modules for motion control: the Directional Motion Control (DMC) module and the Motion Intensity Modulator (MIM) module. These modules work together to allow controlled motion in terms of direction and intensity, as well as to coordinate interaction with the generated video content. We first introduce the test-time, training-free direction control, followed by details on the intensity encoding process. And then we will introduce the motion intensity modulator to modulate the motion intensity and generate videos with the given strength.
@@ 4.1 Training-Free Directional Motion Control
Algorithm 1 Training-Free Directional Motion Control
1:Input text $y$, object phrase $y_{n}$, user-defined trajectory $\mathcal{T}$, denoiser $D$, guidance strength $\eta$, temporal smoothness weight $\lambda$, final guidance timestep $t_{g}$
2:Initialize latent variable $\boldsymbol{z}^{(T)}\sim\mathcal{N}(0,I)$
3:for $t=T$ to $t_{g}$ do
4:     $\sigma_{t}=\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}$# Compute noise level
5:     $A_{i}^{(\gamma)}=\text{Attention}(y_{n},u,i)$# Compute attention
6:     $B_{i}=\text{Region}(\mathcal{T},i)$
7:     $E_{i}\left(A_{i}^{(\gamma)},B_{i},n\right)=\left(1-\frac{\sum_{u\in B_{i}}A_{i,u,n}^{(\gamma)}}{\sum_{u}A_{i,u,n}^{(\gamma)}}\right)^{2}$# Compute energy function for frame $i$
8:     $\mathcal{L}_{s}=\mathbb{E}_{i}\left[\|A_{i,:,:}^{(\gamma)}-A_{i-1,:,:}^{(\gamma)}\|^{2}\right]$# Compute temporal smoothness loss
9:     $G=\nabla_{\boldsymbol{z}^{(t)}}\sum_{i=1}^{L}\sum_{\gamma\in\Gamma}\left(E_{i}\left(A_{i}^{(\gamma)},B_{i},n\right)+\lambda\mathcal{L}_{s}\right)$# Compute gradient
10:     $\boldsymbol{z}^{(t)}\leftarrow\boldsymbol{z}^{(t)}-\sigma_{t}^{2}\eta G$# Update latent variable
11:end for
12:return Generated video from $\boldsymbol{z}^{(0)}$
Cross-attention layers are essential for managing spatial layouts in generated content [15, 11, 7]. The attention score $A_{u,n}^{(\gamma)}$ between spatial location $u$, text token $y_{n}$ in cross-attention layer $\gamma$ determines their association, with $\sum_{n=1}^{N}A_{u,n}^{(\gamma)}=1$. This competitive interaction among text tokens helps guide object motion by biasing attention maps. By aligning attention focus along the trajectory $\mathcal{T}$ with the token $y_{n}$, we gain fine-grained frame-by-frame control over object placement without requiring additional training.
@@
where $A_{i,u,n}^{(\gamma)}$ denotes the attention score at video timestep $i$ in layer $\gamma$, spatial location $u$, and text token $y_{n}$. This function encourages attention to concentrate within the bounding box $B_{i}$ at each timestep $i$, thereby achieving effective alignment of the attention with the trajectory throughout the video sequence.
Temporal Smoothness Loss.
To ensure temporal coherence and avoid abrupt changes in motion, we introduce a temporal smoothness loss across frames, defined as the expectation of squared differences in attention maps over consecutive frames:
$\mathcal{L}_{s}=\mathbb{E}_{i}\left[\|A_{i,:,:}^{(\gamma)}-A_{i-1,:,:}^{(\gamma)}\|^{2}\right],$ (5)
where $A_{i,:,:}^{(\gamma)}$ and $A_{i-1,:,:}^{(\gamma)}$ are attention maps for consecutive frames $i$ and $i-1$. This loss penalizes large frame-to-frame attention changes, promoting smooth motion across frames.
@@
$\boldsymbol{z}^{(t)}\leftarrow\boldsymbol{z}^{(t)}-\sigma_{t}^{2}\eta\nabla_{\boldsymbol{z}^{(t)}}\sum_{i=1}^{L}\sum_{\gamma\in\Gamma}\left(E_{i}\left(A_{i}^{(\gamma)},B_{i},n\right)+\lambda\mathcal{L}_{s}\right),$ (6)
where $\eta>0$ controls the guidance strength, $\lambda$ controls the scale of the temporal smoothness loss, and $\sigma_{t}=\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}$ adjusts for the noise level at diffusion timestep $t$. Scaling by $\sigma_{t}^{2}$ ensures that early diffusion timesteps with high noise receive weaker guidance, which intensifies as noise decreases in later timesteps, refining latent representations near the end of the diffusion process.
To generate a video, we alternate between gradient updates and denoising steps, iteratively refining the latent variable $\boldsymbol{z}^{(t)}$ to align cross-attention maps with the intended trajectory. The combination of iterative guidance and smoothness enforcement allows the model to produce coherent video sequences with precise spatial control. The overall algorithm of the DMC module is shown in Algorithm 1.
4.2 Motion Intensity Modulator Module
The Motion Intensity Modulator (MIM) module enables controlled adjustment of motion intensity in generated videos. During training, we compute the optical flow [16] strength between consecutive frames to quantify motion intensity, using this information as a conditioning input for the diffusion model. This conditioning allows the model to capture varying levels of motion intensity, guiding the generation process based on the specified motion dynamics.
Optical Flow Map Generation.
To extract motion intensity from the video, we first generate optical flow maps that visualize the motion dynamics across frames. Optical flow maps provide pixel-wise information on both the direction and magnitude of movement between consecutive frames, enabling a quantifiable representation of motion intensity. For each video, we sequentially read frames and convert them to grayscale, which allows efficient optical flow calculation using the Farneback method [10]. This dense optical flow method computes the motion between frames. From these components, we derive the magnitude, and the magnitude values provide a measure of motion intensity.
@@ 4.2 Motion Intensity Modulator Module
To incorporate the motion intensity into our model, we convert the optical flow strength of the video into motion embeddings. Similar to [22], rather than feeding the motion intensity as a raw numerical value, we transform it into a word form before training. This aligns with the CLIP model’s preference for text-based representations, as CLIP serves as our text encoder. This approach leverages CLIP’s capability to interpret motion intensity as a semantic feature.

### v2

@@ Abstract
Recent advancements in diffusion models have shown great promise in producing high-quality video content. However, efficiently training video diffusion models capable of integrating directional guidance and controllable motion intensity remains a challenging and under-explored area. To tackle these challenges, this paper introduces Mojito, a diffusion model that incorporates both motion trajectory and intensity control for text to video generation. Specifically, Mojito features a Directional Motion Control (DMC) module that leverages cross-attention to efficiently direct the generated object’s motion without training, alongside a Motion Intensity Modulator (MIM) that uses optical flow maps generated from videos to guide varying levels of motion intensity. Extensive experiments demonstrate Mojito’s effectiveness in achieving precise trajectory and intensity control with high computational efficiency, generating motion patterns that closely match specified directions and intensities, providing realistic dynamics that align well with natural motion in real-world scenarios. Project page: https://sites.google.com/view/mojito-video.
Figure 1: Mojito generates videos that accurately follow specified directions, locations, and trajectories, while adapting to varying input motion intensities. (a) Directional Motion Control: the object (leaf, firefly) in the generated videos can follow input bounding boxes or trajectories over time. (b) Motion Intensity Modulator: increasing input motion intensity levels results in a corresponding increase in motion, transforming a relatively static scene into one with more dynamic movement. Additional examples can be found at https://sites.google.com/view/mojito-video.
@@ 1 Introduction
“Movement is life; control brings mastery.”
—Adapted from Aristotle’s philosophy
Figure 2: Overview of the Mojito framework. In the training pipeline (top),  Mojito uses a VAE Encoder to transform input frames into latent features, processed by Spatial and Temporal Transformers within the U-Net. Motion intensity control is introduced through the Motion Intensity Modulator, consisting of the Optical Flow Map Generator and the Motion Intensity Encoder. The Directional Motion Control module interprets object phrases within the prompt to align attention with specified trajectories. During inference (bottom), Mojito generates videos following user-defined motion intensity and directional guidance.
Text-to-video (T2V) generation (Blattmann et al., 2023a; He et al., 2022; Zheng et al., 2024; Lab & etc., 2024; Xu et al., 2024; Yang et al., 2024; Vchitect, 2024) aims to produce diverse, high-quality videos with given text prompts. Unlike image generation (Rombach et al., 2022; Saharia et al., 2022b; Nichol et al., 2021; Ramesh et al., 2021; Yu et al., 2022; Chang et al., 2023), which produces a single static frame, video generation can extend beyond visual synthesis and has the potential to serve as a world modeling tool (NVIDIA et al., 2025; Xiang et al., 2024; Wang et al., 2024a; Kang et al., 2024) and a physical AI engine (Authors, 2024) for simulating real-world dynamics. As world modeling expects video generation models to control and generate the motion of physical objects, enabling realistic interactions with the environment, existing T2V models—despite their creative and powerful generative capabilities—struggle to produce large motion magnitudes (Zheng et al., 2024; Lab & etc., 2024; Pika, 2024; Wang et al., 2023a) or adaptable motion intensities (i.e., motion speed and magnitude) that align with user intent (NVIDIA et al., 2025; Chen et al., 2024a), limiting their applicability.
@@
Towards this, in this work, we aim to address a new research question: Can text-to-video models be trained efficiently to control both motion direction and intensity in alignment with user intent? The primary limitations of current models stem from several key challenges. First, existing diffusion-based T2V models—whether U-Net or DiT-based (Peebles & Xie, 2022)—rely on diffusion processes in either pixel space or latent visual feature space during training, but lack dedicated motion modeling mechanisms. Consequently, they inherently struggle to adjust object motion in generated videos solely through text prompts and often fail to maintain temporal consistency in moving objects, resulting in flickering, trajectory inconsistencies, and visual artifacts across frames; Second, capturing relative motion data is inherently complex in real-world videos due to the simultaneous movements of both cameras and objects, leading to lack of video data in training. There is also a lack of large-scale, annotated datasets specifically for motion direction and intensity. Existing video datasets rarely include detailed labels for motion dynamics, and obtaining human annotations is labor-intensive; Third, training T2V models with such detailed annotations would also bring substantial computational resources.
@@
To tackle these challenges, we present Mojito, the first diffusion model for text-to-video generation that simultaneously integrates text-based prompts with flexible motion controls, allowing for the precise modulation of both motion direction and intensity. Figure 1 showcases how Mojito enables precise control over both motion trajectory and intensity. To achieve this, Mojito introduces two novel modules apart of our trained video generation backbone: (1) the Motion Intensity Modulator (MIM), which encodes input motion intensities as features and integrates them seamlessly into the diffusion framework during training. (2) the Directional Motion Control (DMC) module, which adjusts object motion direction through cross-attention guidance without training, aligning object trajectories with specified paths.
@@
  • We propose Mojito, a novel diffusion-based transformer framework for text-to-video generation, integrating both trajectory/direction-aware motion direction control and motion intensity modulation within a unified architecture.
  • We introduce a training-free, test-time Directional Motion Control (DMC) method that grants text-to-video generation with user-specified motion trajectories or directions. Our study explores how to effectively guide attention during video generation by analyzing attention maps and identifying the optimal strategy for steering motion using user-defined inputs, such as bounding boxes, during the forward diffusion process. Furthermore, we show that DMC is a model-agnostic approach that generalizes across different video generation architectures, including U-Net-based and DiT-based (Peebles & Xie, 2022) models.
  • To tackle object inconsistency in generated videos under directional motion control, we introduce a novel temporal smoothness function, optimizing it as it updates latent features during the forward diffusion process to enhance coherence across frames.
  • We develop a training-based Motion Intensity Modulator (MIM) that seeks to achieve precise control over motion intensity by leveraging optical flow (Horn & Schunck, 1981), ensuring generated videos align with user-defined motion intensity levels. Additionally, we explore three alternative designs for MIM. We find that combining motion intensity embeddings with text embeddings yields the best performance
  • We demonstrate that Mojito not only generates high-quality videos comparable to state-of-the-art models but also provides effective and efficient motion control. Through ablations, we provide insights into designing efficient, motion-controllable video generation techniques, shedding light on advancing future motion-augmented video generative models.
@@ 2.1 Diffusion-based Text to Video Generation
With the rapid development of generative models, particularly diffusion models, numerous breakthroughs have been achieved in fields such as image generation (Rombach et al., 2022) and video generation (He et al., 2022; Xu et al., 2024; Zheng et al., 2024; Lab & etc., 2024; Chen et al., 2024a; Yang et al., 2024; Vchitect, 2024). Regarding text-to-video (T2V) models, Make-A-Video (Singer et al., 2022), Imagen Video (Saharia et al., 2022a) are cascaded models, while most other works, such as LVDM (He et al., 2022), ModelScope (Wang et al., 2023b), and Align your Latents (Blattmann et al., 2023b), are Stable Diffusion-based models. They extend the SD framework to videos by incorporating temporal layers to ensure temporal consistency among frames and the spatial parameters are inherited from the pretrained SD UNet.  Mojito follows a similar architectural approach, using spatial transformers for spatial information and accepting directional control, and incorporating temporal transformers to handle temporal coherence.
@@ 2.2 Controllable Text-to-Visual Generation
Recent advancements in text-to-visual generation have enabled more precise and stable control over generated outputs. In image generation, several models (Mou et al., 2023; Huang et al., 2023; Zhao et al., 2023; Qin et al., 2023) have been developed to enhance control by leveraging semantic guidance from text prompts alongside various structural conditions. Some approaches utilize attention maps and bounding boxes to manage image layouts, enabling generation based on regional specifications (Chefer et al., 2023; Yang et al., 2023; Li et al., 2023; Chen et al., 2023; Feng et al., 2022).  Mo et al. (2024) presents a training-free method for controllable image generation; In video generation, which involves additional complexity due to motion, recent works have explored methods for motion control. Training-based methods, such as those in Zhang et al. (2024); Wang et al. (2024b), allow for trajectory-based motion control, while Jain et al. (2024) achieves video control through designed masking in the diffusion process. Unlike these approaches, our method leverages attention maps in a training-free manner for directional motion control. Our model can control the movement of specific objects within the generated video, and uniquely, it enables motion intensity control—providing a new level of flexibility in video generation not addressed in previous works.
@@ 3 Preliminaries
Latent Diffusion Models (LDMs), such as Stable Diffusion (Rombach et al., 2022), consist of two main components: an autoencoder and a denoising network. The autoencoder encodes images into a latent space, while the denoising network operates within this latent space, progressively transforming random noise into a coherent image representation. Building on this, Latent Video Diffusion Models (LVDMs) (He et al., 2022) extend LDMs to video generation by incorporating temporal dynamics in the latent space. They are trained using a noise-prediction objective, minimizing the difference between predicted and actual noise added to the data:
@@
4 The Mojito Video Generation Framework
4.1 Methodology Overview
Figure 2 illustrates the architecture of Mojito, which comprises two core modules for motion control: the Directional Motion Control (DMC) module and the Motion Intensity Modulator (MIM). These two modules work together to allow controlled motion in terms of direction and intensity, as well as to coordinate interaction with the generated video content. We first introduce the test-time, training-free  Directional Motion Control. Then we will introduce the Motion Intensity Modulator and finally the Mojito backbone.
@@ 4.2 Training-Free Directional Motion Control
Cross-attention layers are essential for managing spatial layouts in generated content (Hertz et al., 2022; Feng et al., 2022; Chen et al., 2024b). The attention score $A_{u,n}^{(\gamma)}$ between spatial location $u$, text token $y_{n}$ in cross-attention layer $\gamma$ determines their association, with $\sum_{n=1}^{N}A_{u,n}^{(\gamma)}=1$. This competitive interaction among text tokens helps guide object motion by biasing attention maps. By aligning attention focus along the trajectory $\mathcal{T}$ with the token $y_{n}$, we gain fine-grained frame-by-frame control over object placement without requiring additional training.
@@
where $A_{i,u,n}^{(\gamma)}$ denotes the attention score at video timestep $i$ in layer $\gamma$, spatial location $u$, and text token $y_{n}$ for the input object phrase. This function encourages attention to concentrate within the bounding box $B_{i}$ at each timestep $i$, thereby achieving effective alignment of the attention with the trajectory throughout the video sequence.
Temporal Smoothness Function.
To ensure temporal coherence and avoid abrupt changes in motion, we introduce a temporal smoothness function across frames, defined as the expectation of squared differences in attention maps over consecutive frames:
$\mathcal{T}_{s}=\mathbb{E}_{i}\left[\|A_{i,:,:}^{(\gamma)}-A_{i-1,:,:}^{(\gamma)}\|^{2}\right],$ (5)
where $A_{i,:,:}^{(\gamma)}$ and $A_{i-1,:,:}^{(\gamma)}$ are attention maps for consecutive frames $i$ and $i-1$. Optimizing this function penalizes large frame-to-frame attention changes, promoting smooth motion across frames.
@@
$\boldsymbol{z}^{(t)}\leftarrow\boldsymbol{z}^{(t)}-\sigma_{t}^{2}\eta\nabla_{\boldsymbol{z}^{(t)}}\sum_{i=1}^{L}\sum_{\gamma\in\Gamma}\left(E_{i}\left(A_{i}^{(\gamma)},B_{i},n\right)+\lambda\mathcal{T}_{s}\right),$ (6)
where $\eta>0$ controls the guidance strength, $\lambda$ controls the scale of the temporal smoothness function, and $\sigma_{t}=\sqrt{\frac{1-\alpha_{t}}{\alpha_{t}}}$ adjusts for the noise level at diffusion timestep $t$. Scaling by $\sigma_{t}^{2}$ ensures that early diffusion timesteps with high noise receive weaker guidance, which intensifies as noise decreases in later timesteps, refining latent representations near the end of the diffusion process.
To generate a video, we alternate between gradient updates and denoising steps, iteratively refining the latent variable $\boldsymbol{z}^{(t)}$ to align cross-attention maps with the intended trajectory. The combination of iterative guidance and smoothness enforcement allows the model to produce coherent video sequences with precise spatial control. The overall algorithm framework of the DMC module is showcased in Algo. 1.
4.3 Motion Intensity Modulator
The Motion Intensity Modulator (MIM) enables controlled adjustment of motion intensity in generated videos. During training, we compute the optical flow (Horn & Schunck, 1981) strength between consecutive frames to quantify motion intensity, using it as a conditioning input for the diffusion model. This conditioning allows the model to capture varying levels of motion intensity, guiding the generation process based on the specified motion dynamics.
Optical Flow Map Generator.
To quantify motion intensity in videos, we compute optical flow maps that capture the temporal dynamics between consecutive frames. These maps encode pixel-wise motion direction and magnitude, providing a structured representation of movement. For each video, frames are processed sequentially and converted to grayscale to facilitate efficient computation using the Farneback method (Farnebäck, 2003). This approach estimates motion by analyzing pixel displacements between adjacent frames. The resulting magnitude values serve as a measure of motion intensity, which we normalize and discretize into integer levels ranging from $1$ to $10$.
@@ 4.3 Motion Intensity Modulator
To incorporate the motion intensity into our model, we convert the optical flow strength of videos into embeddings. Similar to Kandala et al. (2024), rather than feeding the motion intensity as a raw numerical value, we transform it into a word form during training. This aligns with the CLIP model’s preference for text-based representations, as CLIP serves as our text encoder.